OpenStudy (anonymous):

Factor 4x^2 – 81y^2.

OpenStudy (anonymous):

keep this identity in your mind \[\huge (a-b)(a+b)=a^2-b^2\] now let \(a=2x\) and \(b=9y\) to get ur answer

OpenStudy (anonymous):

you might need to help me a bit further i dont quite understand

OpenStudy (anonymous):

@mukushla please

OpenStudy (anonymous):

note that \(4x^2 – 81y^2=(2x)^2-(9y)^2\)

OpenStudy (anonymous):

thank u!

OpenStudy (anonymous):

yw :)